OpenStudy (anonymous):

What polynomial has roots of -6, -4, and 1 ? x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

OpenStudy (cwrw238):

(x+6)(x+4)(x-1) = 0 has these roots which will expand to one of your options

OpenStudy (cwrw238):

= (x-1)(x^2 +10+24)

OpenStudy (cwrw238):

that narrows it down to 2nd or 4th (-24)

OpenStudy (cwrw238):

that should be 10x in second parentheses

OpenStudy (cwrw238):

expansion will give -10x +24x - so which ne do you think it is?

OpenStudy (anonymous):

i think the 3rd one?

OpenStudy (cwrw238):

nope -10x + 24x = +14x

OpenStudy (cwrw238):

check out the one with +14x in it

OpenStudy (cwrw238):

it cant be the 3rd one as it contains +24 - - see above - it must contain -24

OpenStudy (anonymous):

shoot ok let me see

OpenStudy (anonymous):

the last one right? i didnt even see the 14x haha

OpenStudy (cwrw238):

right - hope you dont mind me 'playing' with you - they like us to guide you through on this site - not just give you the answer

OpenStudy (anonymous):

no i understand i rather understand how to solve it then just get an answer :]